VOCM (a radio station) trasmits at 590 AM. This means that the radio waves carrying the signal have a frequency of 590 kHz. Given this, how long does it take for 1 complete cycle of the radio wave to leave the station?
Answer #1

Since we know that T = 1/f

T = 1/(590000) = 1.69 X 10-6 S